# Modular degree and a conjecture of Watkins

### v2

@@ 4. On the growth of modular degree
In this section, we shall discuss the growth properties of $m_{E}$ in both $\mathbb{Q}$ and functional field situations. Let us first consider $E/\mathbb{Q}$ to be an elliptic curve of conductor $N$. It is conjectured that $m_{E}=O(N^{2+\varepsilon})$. Now recall the two identities from (1). It follows from the proof of [29, Theorem 1] that $N^{1-\varepsilon}<||f_{E}||^{2}_{N}<N^{1+\varepsilon}$. From [41, Lemma 2.1] we have an explicit lower bound of the form $m_{E}>\frac{60\pi^{2}N^{1-\varepsilon}}{D_{E}^{1/6}}$. Watkins [41] showed that $L(\mathrm{Sym}^{2}(E),1)\gg\frac{1}{N^{(2)}}$, by showing that there exists no real zero for $L(\mathrm{Sym}^{2}(E),s)$ in a region of the form $\mathrm{Re}(s)\geq 1-\frac{C}{\log N^{(2)}}$, where $N^{(2)}$ is conductor of the associated L-function, which is $\geq N_{E}^{2}$. This gives a lower bound on $m_{E}$ of order $\gg N^{7/6-\varepsilon}$. Here $L(\mathrm{Sym}^{2}(E),s)$ is the L-function associated with the motive $H^{1}(\mathrm{Sym}^{2}(E))$, whose local factors $L_{p}(s)$, are of the polynomials in $p^{-s}$, that are determined by $\mathrm{det}(1-\mathrm{Frob}_{p}|{\mathrm{Sym}^{2}}(T_{\ell}(E)\otimes\mathbb{Q}_{\ell}))$, for any prime $p\neq\ell$. The existence of the zero-free region essentially shows that $m_{E}\gg\frac{N^{7/6}}{\log N}\prod_{p^{\mid}N}L_{p}(1)$. The local factors of the (motivic) Symmetric square $L$-function is easier to understand at the primes of good and multiplicative reductions. For the other primes, these factors are $\geq 1$ (see [41, page 5, Corollary]), whenever $p\equiv 1\pmod{12}$, and hence the lower bound follows.
One approach to get an upper bound would be to know the bounds for Manin’s constant and Falting’s height of $E$. Indeed, for any integer $A,B$, consider the Frey curve $E:y^{2}=x(x-A)(x+B)$. It turns out that the Frey curves are semistable away from $2$. In particular, $c_{E}$ is uniformly over this family. Then a computation of $j$ and $\gamma$ of $E_{A,B,C}$, gives the connection with the ABC problem plugging them in (1). In particular, it turns out that ABC implies the conjectural degree bound holds. In fact, any unconditional result on ABC contributes an unconditional bound on the modular degree. For instance, [38] implies a bound of order $e^{O_{\varepsilon}(N^{1/3+\varepsilon})}$.
Murty in [29] considered the conjecture for elliptic curves with CM. Other than the ones with $j$ invariants, respectively 00 and $1728$, it turns out that any elliptic curve with CM is the quadratic twist of a set of finitely many elliptic curves $\mathcal{E}$. Note that $j_{E}=j_{E^{(D)}}$ for any $E\in\mathcal{E}$, and in particular, $h(E)-h(E^{(D)})=\frac{1}{2}\log D$. This shows that
(16) $(1-\varepsilon)\log{N_{D}}+\log D\ll_{E}\log m_{E^{(D)}}\ll_{E}(1+\varepsilon)\log{N_{D}}+\log D,$
@@
For any set of finitely many primes $S$, denote $\mathcal{E}_{S}$ be the set of all elliptic curves over $E/\mathbb{Q}$ for which $N_{E}$ is semistable away from $S$. Note that $c_{E}=O_{S}(1)$ uniformly over this family. It is known due to Fouvry et al. [16] that $\log(\Delta_{E})\leq(1+\varepsilon)\log N_{E}$ for almost all elliptic curves over $E/\mathbb{Q}$, when arranged by height. On the other hand for any $\delta>0$, the number of pairs $(A,B)\in\mathbb{Z}^{2}$, with $\left|\frac{A^{3}}{\Delta_{E}}\right|<1+\delta$ has density at most $\delta$. Choosing $\delta$ suitably, we now have that $h(j_{E})=\log(\Delta_{E})\leq(1+\varepsilon)\log N_{E}$, for all elliptic curves outside a set of density $O(\varepsilon)$. We now claim that, $\gamma=O_{S}(1)$, for any $E\in\mathcal{S}$. Modulo this, we have $m_{E}\ll_{S}N_{E}^{7/6+\varepsilon}$, for any $E\in\mathcal{E}_{S}$. It is then enough to prove the claim and show that $\mathcal{E}_{S}$ has a density at least $1-\prod_{p\in S}\frac{1}{p^{2}}$.
First, to prove the claim, note that $E$ has an additive reduction for any prime $p$ dividing $\gamma$. Since $E\in\mathcal{S}_{p}$, it is evident that $p\in S$. It is enough to bound the power of $p$ by dividing $\gamma$. For this, we need to chase Kodaira’s classification. For instance, from Table 1 in [36], we immediately see that the power is bounded for all the types, other than $I_{n},I^{*}_{n}~{}(n\geq 1)$. Following the notations of [36, Theorem 1.6], note $j_{E}=\frac{16a^{2}-48b}{\Delta}$, and the $p$-adic valuation of numerator of at most $6$, for any prime $p$ and for any $n\geq 2$. This shows that the valuation of $\Delta_{E}$ in $j_{E}$ does not go down too much, and in particular, this analysis shows that $\nu_{p}(\gamma)\leq 6$ for any prime $p\in S$.
Now to compute the density of $\mathcal{E}_{S}$, again note from Table 1 in [36] that, the proportion of all elliptic curves having additive reduction at a prime $p$ is $1/p^{2}$. This shows that, the density of $\mathcal{E}_{S}$ is at least $01-\left(\sum_{p\not\in S}\frac{1}{p^{2}}\right)>\frac{1}{2}+\sum_{p\in S}\frac{1}{p^{2}}>$ 0, as desired.
Now let $E/\mathbb{Q}$ be any given elliptic curve, and $\{E^{(D)}\}$ be the family of quadratic twists. For any $D$ co-prime to $N_{E}$, we have $N_{E^{(D)}}=ND^{2}$. Then it follows from (16) that
@@
Remark 4.
It follows from the proof of Theorem 11 that a stronger estimate $m_{E}\ll N_{E}^{7/6+\varepsilon}$. Assuming Manin’s conjecture, i.e., that Manin’s constant is uniformly bounded, the same estimate with exponent $7/6+\varepsilon$ holds for almost all elliptic curves. Moreover, the reader also may note from the proof of Theorem 11 that exponent $3/2+\varepsilon$ can be achieved for a positive proportion of elliptic curves in any family of quadratic twists.
Note that the degree bound could also be seen as one of the consequences for the bound of $||f_{E}||^{2}$. Let us now quickly mention another consequence. Rankin in [33, Theorem 1] showed that $\sum_{1\leq n\leq x}|a_{E}(n)|^{2}=\alpha_{E}x^{2}+o(x^{2}),$ where $\alpha_{E}>0$ is some computable constant. In particular, applying partial summation formula, it follows from [29] that $\alpha_{E}=O\left(\frac{\phi(N)||f_{E}||^{2}}{N^{2}}\right)=O(\log N)$. We would also like to point out that an inexplicit version of the square-sum problem for any Fuchsian group (of the first kind) is discussed in [4].
Regarding the degree bound conjecture, one may also ask for a weaker problem; is any prime factor $\ell_{E}$ of $m_{E}$ is $O(N^{2+\varepsilon})$? Now one may ask about the growth of $\ell_{E}$, varying over all the elliptic curves $E/\mathbb{Q}$. Note that any such $\ell_{E}$ is also a congruence prime for $E$, i.e., there exists a cuspform $g(z)=\sum_{n>0}a_{g}(n)q^{n}$ with integer Fourier coefficients $b_{n}$, such that $a_{E}(n)\equiv a_{g}(n)\pmod{\ell_{E}}$ for all $n$. Suppose that $g$ corresponds to an elliptic curve $E^{\prime}/\mathbb{Q}$. Then the representation $\rho_{E,E^{\prime}}:\mathrm{Gal}_{\mathbb{Q}}\to\mathrm{GL}_{2}(\mathbb{Z}/\ell_{E}\mathbb{Z})$ is not subjective. With this observation, we can prove the following.
Theorem 12.
Let $E/\mathbb{Q}$ be any elliptic curve, and $\ell_{E}\geq 5$ be a prime diving $m_{E}$. Set $\mathcal{E}(x)$ be the set of all elliptic curves over $\mathbb{Q}$ of height at most $x$. Then, $g$ corresponds to at most $O(x^{4}\log^{2}x)$ many elliptic curves in $\mathcal{E}(x)$.
To prove this, we follow the similar Sieve approach as in the proof of (10) in [23]. The difference is that, we have to carry out the sieving on copy of $\mathbb{Z}^{2}$ inside $\mathbb{Z}^{4}$.
Proof of Theorem 12.
@@
where $\pi_{2}$ denotes the natural projection onto the second component. Regardless of what the first component $E$ is, $\Omega_{E}(C)$ is in bijection with $\{(r,s)\in(\mathbb{Z}/p\mathbb{Z})^{2}:\pi_{2}(\rho_{E_{r,s}}(\mathrm{Frob}(p)))\in C\}$, provided that $p=1\pmod{\ell_{E}}$. In that case, we have from [22, Theorem 8] that #$\#\Omega_{E}(C)=p^{2}\left(\frac{\#C}{\#\mathrm{SL}_{2}(\mathbb{Z}/\ell_{E}\mathbb{Z})}+O\left(\frac{\#C}{p^{1/2}}\right)\right)$. This shows that,
@@
where $\pi(x,\ell_{E},1)$ denotes $\{p\leq x:p=1\pmod{\ell_{E}}\}$. Let us now denote $\mathcal{E}_{E}(x)$ be the set of all elliptic curves $E^{\prime}$ in $\mathcal{E}(x)$, that corresponds to $g$. Since $\mathrm{im}(\rho_{E,E^{\prime}})\neq\Delta(\ell_{E})$ for any such $E^{\prime}$, by [23, Lemma 3.1] we know that, there exists conjugacy classes $C_{1},C_{2}$ in $\mathrm{SL}_{2}(\mathbb{Z}/\ell_{E}\mathbb{Z})$ such that $\mathrm{im}(\rho_{E,E^{\prime}})\cap C_{1}\times C_{2}=\phi$. By large sieve, we argue similarly as in [23, page 3389], and get for any conjugacy class $C$ that
$\pi(x,\ell_{E},1)^{2}\#\mathcal{E}^{2}_{E}(x)\leq\sum_{E^{\prime}\in\mathcal{E}(x)}(\pi_{E,E^{\prime}}(x,C)-\delta_{C}\pi(x,\ell_{E},1))^{2}\ll\#\mathcal{E}(x)P_{C}(x),$
where $\delta_{C}=\frac{\#C}{\#\mathrm{SL}_{2}(\mathbb{Z}/\ell_{E}\mathbb{Z})}$, and $\pi_{E,E^{\prime}}(x,C)=\{p\leq x:\pi_{2}(\rho_{E,E^{\prime}}(\mathrm{Frob}_{p}))\in C\}$. In particular, #$\#\mathcal{E}^{2}_{E}(x)\ll\sum_{C_{1}\times C_{2}}\frac{\#\mathcal{E}(x)P_{C}(x)}{\pi(x,\ell_{E},1)^{2}}\ll_{\ell_{E}}x^{4}(\log x)^{2}$, as desired. ∎
Remark 5.
Of course, $g$ is not necessarily an eigenform. Then we can write, $g$ as linear combination of a family of hecke eigenforms $\{f_{i}\}_{s(N)}$, where $s(N)=\mathrm{dim}(S_{2}(\Gamma_{0}(N)))$. We can then consider the family of representations $\{\rho_{E,E_{1},E_{2},\cdots,E_{s_{2}(N)}}\}_{\begin{subarray}{c}E_{i}\in\mathcal{E}(x)\\ 1\leq i\leq s(N)\end{subarray}}$, and draw a similar conclusion as in Theorem 12, applying the multi dimensional analog of Large sieve.
@@ 4.1. Function field analog of growth
Following the same set us as in Section 3.7, we denote $|\mathfrak{n}|_{\infty}=q^{\mathrm{deg}(\mathfrak{n})}$, and $\mathrm{deg}_{\mathrm{ns}}(j_{E})$ be the nonseparable degree of $\mathbb{F}_{q}(T)/\mathbb{F}_{q}(j_{E})$. Given any newform $f_{E}$, it turns out that $L(Sym^{2}f_{E},2)=q\frac{||f_{E}||^{2}}{\mathfrak{n}}$. Moreover, $\mathfrak{m}_{E}=\frac{||f_{E}||^{2}}{\mathrm{val}_{\infty}(j_{E})}$, where $\mathrm{val}_{\infty}(j_{E})$ denotes the number of irreducible components of the singular fiber $\infty$ (note that, we always assume that $E$ has bad reduction at $\infty$). Of course $1\leq\mathrm{val}_{\infty}(j_{E})\leq\mathrm{deg}(\Delta_{E})$, and hence, $\frac{||f_{E}||^{2}}{\mathrm{deg}(\Delta_{E})}\leq\mathfrak{m}_{E}\leq||f_{E}||^{2}$. Moreover, it follows from [32] that $\mathrm{deg}(\Delta_{E})\leq\mathrm{deg}_{\mathrm{ns}}(j_{E})\mathrm{deg}(\mathfrak{n}_{E})$, where $\mathrm{deg}_{\mathrm{ns}}(j_{E})$ denotes the non-seperable degree of $\mathbb{F}_{q}(T)/\mathbb{F}_{q}(j_{E})$. Papikian [30] provided both analytic and algebraic ways to bound the norm of $f_{E}$. In the analytic approach, the bound follows as a consequence of Ramanujan’s conjecture on the holomorphic $L$-functions on $\mathbb{C}$, and Rademacher’s version of Phragmen-Lindelof’s theorem for the estimation of $L$-functions on a strip. Now for the algebraic approach, Grothendieck’s theory L-function implies that $L(s,\mathrm{Sym}^{2}(E))$ can be written as a quotient of two polynomials in $q^{-s}$, whose degree is bounded by $42\mathrm{deg}(\mathfrak{n}_{E})-$ 4. This is a consequence of Grothendieck–Ogg–Shafarevich’s formula, realizing $\mathrm{Sym}^{2}$ as a constructible sheaf over $\mathbb{P}^{1}_{\mathbb{F}_{q}}$. To get the degree bound, they assumed that $E$ is semi-stable, which gurantees that all the wild parts $\delta_{\mathfrak{p}}|\mathfrak{p}\in\mathbb{P}^{1}_{\mathbb{F}_{q}}$ are 00. However, a weaker bound of the form $\ll\mathrm{deg}(\mathfrak{n}_{E})$ (on the degree of $L(s,\mathrm{Sym}^{2}(E))$), saves the day for us. For this, we may put a weaker (than being semistable) condition on $E$, that sum of all the wild parts, i.e., $\sum_{\mathfrak{p}\in\mathbb{P}^{1}_{\mathbb{F}_{q}}}\delta_{\mathfrak{p}}$ is $O(\mathfrak{n}_{E})$. Conversely, a lower bound is also achieved using some standard analytic tools. To summarize the discussion, we now can state the bounds [30, Corollary 6.2] in a nice geometrical way;
@@
Now the question that naturally comes to our mind is that, whether it is possible to remove the extra dependency on $E$, that appears as a factor $\mathrm{deg}_{\mathrm{ns}}(j_{E})$ in the lower bound. As pointed out by Papikian in [30, page 347] that, on the family $\{\Delta_{E^{q^{n}}}\}_{n\geq 1}\}$, discriminant increases, but the conductor remains same, as $n$ grows. Hence, the extra factor can not be removed for all elliptic curves. Following the spirit as in all the previous sections, we consider the family of all elliptic curves, and quadratic twists, to study the proportion of elliptic curves where it is possible to remove the factor. First, given any elliptic curve $E:=y^{2}=x^{3}+f(T)x+g(T)$ with $f(T),g(T)\in\mathbb{F}_{q}[T]$, we define height $H(E)=\max\left\{3\mathrm{deg}(f),2\mathrm{deg}(g)\right\}$. This is the function field analog of height in the $\mathbb{Q}$ case. Denote $\mathcal{E}_{\mathbb{F}}(d)$ be the number of elliptic curves of height at most $d$, which is precisely $q^{5d/6}$. Now note that
@@
This shows that for any integer $D\leq d$, we have the bound $m_{E}\gg\frac{\mathrm{deg}(\mathfrak{n}_{E})}{D}$, for least $q^{5D/6}$ many $E$ in $\mathcal{E}_{\mathbb{F}}(d)$. On the hand, as long as we consider the families of quadratic twists, we see that $j_{E}$ remains invariant. Therefore, in any such family, the factor $\mathrm{deg}_{\mathrm{ns}}(j_{E})$ remains unchanged.

### v3

@@ 4. On the growth of modular degree
[…]
One approach to get an upper bound would be to know the bounds for Manin’s constant and Falting’s height of $E$. Indeed, for any integer $A,B$, consider the Frey curve $E:y^{2}=x(x-A)(x+B)$. It turns out that the Frey curves are semistable away from $2$. In particular, $c_{E}$ is uniformly over this family. Then a computation of $j$ and $\gamma$ of $E_{A,B,C}$, gives the connection with the ABC problem plugging them in (2.1). In particular, it turns out that ABC implies the conjectural degree bound holds. In fact, any unconditional result on ABC contributes an unconditional bound on the modular degree. For instance, [SY01] implies a bound of order $e^{O_{\varepsilon}(N^{1/3+\varepsilon})}$.
Murty in [Murty99] considered the conjecture for elliptic curves with CM. Other than the ones with $j$ invariants, respectively 0 and $1728$, it turns out that any elliptic curve with CM is the quadratic twist of a set of finitely many elliptic curves $\mathcal{E}$. Note that $j_{E}=j_{E^{(D)}}$ for any $E\in\mathcal{E}$, and in particular, $h(E)-h(E^{(D)})=\frac{1}{2}\log D$. This shows that
(4.1) $(1-\varepsilon)\log{N_{D}}+\log D\ll_{E}\log m_{E^{(D)}}\ll_{E}(1+\varepsilon)\log{N_{D}}+\log D,$
@@
For any set of finitely many primes $S$, denote $\mathcal{E}_{S}$ be the set of all elliptic curves over $E/\mathbb{Q}$ for which $N_{E}$ is semistable away from $S$. Note that $c_{E}=O_{S}(1)$ uniformly over this family. It is known due to Fouvry et al. [FNT92] that $\log(\Delta_{E})\leq(1+\varepsilon)\log N_{E}$ for almost all elliptic curves over $E/\mathbb{Q}$, when arranged by height. On the other hand for any $\delta>0$, the number of pairs $(A,B)\in\mathbb{Z}^{2}$, with $\left|\frac{A^{3}}{\Delta_{E}}\right|<1+\delta$ has density at most $\delta$. Choosing $\delta$ suitably, we now have that $h(j_{E})=\log(\Delta_{E})\leq(1+\varepsilon)\log N_{E}$, for all elliptic curves outside a set of density $O(\varepsilon)$. We now claim that, $\gamma=O_{S}(1)$, for any $E\in\mathcal{E}$. Modulo this, we have $m_{E}\ll_{S}N_{E}^{7/6+\varepsilon}$, for any $E\in\mathcal{E}_{S}$. It is then enough to prove the claim and show that $\mathcal{E}_{S}$ has a density at least $1-\left(\sum_{p\not\in S}\frac{1}{p^{2}}\right)$.
First, to prove the claim, note that $E$ has an additive reduction for any prime $p$ dividing $\gamma$. Since $E\in\mathcal{E}_{p}$, it is evident that $p\in S$. It is enough to bound the power of $p$ by dividing $\gamma$. For this, we need to chase Kodaira’s classification. For instance, from Table 1 in [SSW21], we immediately see that the power is bounded for all the types, other than $I_{n},I^{*}_{n}~(n\geq 1)$. Following the notations of [SSW21, Theorem 1.6], note $j_{E}=\frac{16a^{2}-48b}{\Delta}$, and the $p$-adic valuation of numerator of at most $6$, for any prime $p$ and for any $n\geq 2$. This shows that the valuation of $\Delta_{E}$ in $j_{E}$ does not go down too much, and in particular, this analysis shows that $\nu_{p}(\gamma)\leq 6$ for any prime $p\in S$.
Now to compute the density of $\mathcal{E}_{S}$, again note from Table 1 in [SSW21] that, the proportion of all elliptic curves having additive reduction at a prime $p$ is $1/p^{2}$. This shows that, the density of $\mathcal{E}_{S}$ is at least $1-\left(\sum_{p\not\in S}\frac{1}{p^{2}}\right)>\frac{1}{2}+\sum_{p\in S}\frac{1}{p^{2}}>0$, as desired.
Now let $E/\mathbb{Q}$ be any given elliptic curve, and $\{E^{(D)}\}$ be the family of quadratic twists. For any $D$ co-prime to $N_{E}$, we have $N_{E^{(D)}}=ND^{2}$. Then it follows from (4.1) that
@@
Remark 4.2.
It follows from the proof of Theorem 4.1 that a stronger estimate $m_{E}\ll N_{E}^{7/6+\varepsilon}$. Assuming Manin’s conjecture, i.e., that Manin’s constant is uniformly bounded, the same estimate with exponent $7/6+\varepsilon$ holds for almost all elliptic curves. Moreover, the reader also may note from the proof of Theorem 4.1 that exponent $3/2+\varepsilon$ can be achieved for a positive proportion of elliptic curves in any family of quadratic twists.
[…]
Regarding the degree bound conjecture, one may also ask for a weaker problem; is any prime factor $\ell_{E}$ of $m_{E}$ is $O(N^{2+\varepsilon})$? Now one may ask about the growth of $\ell_{E}$, varying over all the elliptic curves $E/\mathbb{Q}$. Note that any such $\ell_{E}$ is also a congruence prime for $E$, i.e., there exists a cuspform $g(z)=\sum_{n>0}a_{g}(n)q^{n}$ with integer Fourier coefficients $b_{n}$, such that $a_{E}(n)\equiv a_{g}(n)\pmod{\ell_{E}}$ for all $n$. Suppose that $g$ corresponds to an elliptic curve $E^{\prime}/\mathbb{Q}$. By Brauer–Nesbitt theorem, the representation $\rho_{E,E^{\prime}}:\mathrm{Gal}_{\mathbb{Q}}\to\Delta(\ell_{E})$ is not surjective, when $\ell_{E}>2$; see [Jones10] for the definition of $\Delta(\ell_{E})$ and this representation. With this observation, we can prove the following.
Theorem 4.3.
Let $E/\mathbb{Q}$ be any elliptic curve, and $\ell_{E}\geq 5$ be a prime diving $m_{E}$. Denote $\mathcal{E}(x)$ be the set of all elliptic curves over $\mathbb{Q}$ of height at most $x$. Then, $g$ corresponds to at most $O(x^{4}\log^{2}x)$ many elliptic curves in $\mathcal{E}(x)$.
To prove this, we follow the similar Sieving approach as in the proof of (10) in [Jones13]. The difference is that, we have to carry out the sieving on a copy of $\mathbb{Z}^{2}$ inside $\mathbb{Z}^{4}$.
Proof of Theorem 4.3.
@@
where $\pi_{2}$ denotes the natural projection onto the second component. Regardless of what the first component $E$ is, $\Omega_{E}(C)$ coincides with $\{(r,s)\in(\mathbb{Z}/p\mathbb{Z})^{2}:\rho_{E_{r,s}}(\mathrm{Frob}(p))\in C\}$, provided that $p=1\pmod{\ell_{E}}$. In that case, we have from [Jones10, Theorem 8] that $\Omega_{E}(C)=p^{2}\left(\frac{\#C}{\#\mathrm{SL}_{2}(\mathbb{Z}/\ell_{E}\mathbb{Z})}+O\left(\frac{\#C}{p^{1/2}}\right)\right)$. This shows that,
@@
where $\pi(x,\ell_{E},1)$ denotes $\{p\leq x:p=1\pmod{\ell_{E}}\}$. Let us now denote $\mathcal{E}_{E}(x)$ be the set of all elliptic curves $E^{\prime}$ in $\mathcal{E}(x)$, that corresponds to $g$. Due to [duke97, Lemma 5], except for at most $x^{4}\log^{2}x$ many $E^{\prime}$ in $\mathcal{E}(x)$, we can assume that $\mathrm{im}(\rho_{E^{\prime}})=\mathrm{GL}_{2}(\mathbb{Z}/\ell_{E}\mathbb{Z})$. Again by Brauer-Nesbitt, we have $\mathrm{im}(\rho_{E})=\mathrm{GL}_{2}(\mathbb{Z}/\ell_{E}\mathbb{Z})$. Since $\mathrm{im}(\rho_{E,E^{\prime}})\neq\Delta(\ell_{E})$ for any such $E^{\prime}$, by [Jones13, Lemma 3.1] we know that, there exists conjugacy classes $C_{1},C_{2}$ in $\mathrm{SL}_{2}(\mathbb{Z}/\ell_{E}\mathbb{Z})$ such that $\mathrm{im}(\rho_{E,E^{\prime}})\cap C_{1}\times C_{2}=\phi$. By large sieve, we argue similarly as in [Jones13, page 3389], and get for any conjugacy class $C$ that
$\pi(x,\ell_{E},1)^{2}\#\mathcal{E}_{E}(x)\leq\sum_{E^{\prime}\in\mathcal{E}(x)}(\pi_{E,E^{\prime}}(x,C)-\delta_{C}\pi(x,\ell_{E},1))^{2}\ll\#\mathcal{E}(x)P_{C}(x),$
where $\delta_{C}=\frac{\#C}{\#\mathrm{SL}_{2}(\mathbb{Z}/\ell_{E}\mathbb{Z})}$, and $\pi_{E,E^{\prime}}(x,C)=\{p\leq x:\pi_{2}(\rho_{E,E^{\prime}}(\mathrm{Frob}_{p}))\in C\}$. Hence, $\mathcal{E}_{E}(x)\ll\sum_{C_{1}\times C_{2}}\frac{\#\mathcal{E}(x)P_{C}(x)}{\pi(x,\ell_{E},1)^{2}}\ll_{\ell_{E}}x^{4}\log^{2}x$, as desired. ∎
Remark 4.4.
Of course, $g$ is not necessarily an eigenform. Then we can write, $g$ as linear combination of a family of hecke eigenforms $\{f_{i}\}_{s(N)}$, where $s(N)=\mathrm{dim}(S_{2}(\Gamma_{0}(N)))$. We can then consider the family of representations $\{\rho_{E,E_{1},E_{2},\cdots,E_{s_{2}(N)}}\}_{\begin{subarray}{c}E_{i}\in\mathcal{E}(x)\\ 1\leq i\leq s(N)\end{subarray}}$, and draw a similar conclusion as in Theorem 4.3, applying the multi dimensional analog of Large sieve.
@@ 4.1. Function field analog of growth
[…]
@@
Now the question that naturally comes to our mind is that, whether it is possible to remove the extra dependency on $E$, that appears as a factor $\mathrm{deg}_{\mathrm{ns}}(j_{E})$ in the lower bound. As pointed out by Papikian in [PM2002, page 347] that, on the family $\{\Delta_{E^{q^{n}}}\}_{n\geq 1}\}$, discriminant increases, but the conductor remains same, as $n$ grows. Hence, the extra factor can not be removed for all elliptic curves. Following the spirit as in all the previous sections, we consider the family of all elliptic curves, and quadratic twists, to study the proportion of elliptic curves where it is possible to remove the factor. First, given any elliptic curve $E:=y^{2}=x^{3}+f(T)x+g(T)$ with $f(T),g(T)\in\mathbb{F}_{q}[T]$, we define height $H(E)=\max\left\{3\mathrm{deg}(f),2\mathrm{deg}(g)\right\}$. This is the function field analog of height in the $\mathbb{Q}$ case. Denote $\mathcal{E}_{\mathbb{F}}(d)$ be the number of elliptic curves of height at most $d$, which is precisely $q^{5d/6}$. Now note that
@@
This shows that for any integer $D\leq d$, we have the bound $m_{E}\gg\frac{\mathrm{deg}(\mathfrak{n}_{E})}{D}$, for least $q^{5D/6}$ many $E$ in $\mathcal{E}_{\mathbb{F}}(d)$. On the other hand, as long as we consider the families of quadratic twists, we see that $j_{E}$ remains invariant. Therefore, in any such family, the factor $\mathrm{deg}_{\mathrm{ns}}(j_{E})$ remains unchanged.